# Ghost Noise for Regularizing Deep Neural Networks

### v1

@@ Abstract
00footnotetext: Code available at https://github.com/epfml/ghost-noise
@@ 1 Introduction
The use of normalization methods has become widespread in deep learning since the introduction of Batch Normalization [12] (BN) in 2015. For convolutional network training, the batch normalization of a tensor $\bm{X}\in\mathbb{R}^{N\times C\times H\times W}$ with batch size $N$, $C$ channels, height $H$, and width $W$, is given by:
$\hat{{\bm{X}}}=\frac{\bm{X}-{\bm{\mu}}}{\sqrt{{\bm{\sigma}}^{2}+\varepsilon}},\qquad{\bm{\mu}}=\frac{1}{NHW}\sum_{n,h,w}\bm{X}_{n,:,h,w},\quad{\bm{\sigma}}^{2}=\frac{1}{NHW}\sum_{n,h,w}(\bm{X}-{\bm{\mu}})^{2}_{n,:,h,w}$ (1)
where ${\bm{\mu}}\in\mathbb{R}^{1\times C\times 1\times 1}$ is the mean, ${\bm{\sigma}}\in\mathbb{R}^{1\times C\times 1\times 1}$ the variance, and operations are broadcasted and performed elementwise. Following a normalization operation, we typically have an affine transformation ${\bm{Y}}={\bm{\gamma}}\odot\hat{\bm{X}}+{\bm{\beta}}$ with a trainable gain ${\bm{\gamma}}\in\mathbb{R}^{1\times C\times 1\times 1}$ and bias ${\bm{\beta}}\in\mathbb{R}^{1\times C\times 1\times 1}$. The mean and variance are computed over the batch dimension causing the final prediction of a given sample to depend on others in the batch. This cross-sample dependency is undesirable for inference, e.g. if test data arrives in small correlated batches or even one sample at a time. In practice, ${\bm{\mu}},{\bm{\sigma}}$ from Equation 1 are replaced with $\hat{{\bm{\mu}}}$ and $\hat{{\bm{\sigma}}}$ that are exponential moving averages of the ${\bm{\mu}},{\bm{\sigma}}$ used during training. Each sample contributes to its own normalization statistics during training but not inference, causing a discrepancy that can degrade performance. This can be addressed by modifying the deployment network to take the statistics of the current sample into account as done in EvalNorm [19] and Inference Example Weighing [21], at a slightly increased computation cost.
@@
The batch dependency of BN can be detrimental during training under certain circumstances. This is especially true when the available batch size is small or the elements of the batch are correlated [25], where the statistics may deviate significantly from the population statistics. This can lead to a mismatch between the statistics used for normalization during training and those used during inference. A number of alternative methods such as Layer Normalization [1], Instance Normalization [10], Group Normalization [24], Online Normalization [4], Batch Renormalization [11] and Weight Normalization [18] have been proposed to overcome some of these issues. These either normalize across other dimensions of the tensor $\bm{X}$, are applied to the weights instead, or make use of the history over prior batches to augment the effective size of the batch.
@@
The stochasticity induced by the cross-sample dependency of BN can sometimes have a beneficial regularization effect, which is frequently observed but poorly understood. From a geometrical point of view, Keskar et al. [13] empirically found that training with small batches arrives at flatter minima, which they argue contributes to better generalization. Hoffer et al. [8] observed that performing batch normalization over smaller subgroups of the batch could improve generalization. They call these smaller batches ghost batches and this method of explicitly using smaller batches in batch normalization is known as Ghost Batch Normalization (GBN) or sometimes just Ghost Normalization. Summers and Dinneen [21] also observed this effect and recommended the use of GBN. The regularization effect arises from the noise in ${\bm{\mu}},{\bm{\sigma}}$ computed over a batch compared to the corresponding statistics for the full dataset. This noise, which we will refer to as Ghost Noise, increases for smaller batches and may explain the improved generalization with Ghost Batch Normalization. Figure 1 shows an example of how the ghost batch size can affect performance.
In this work we analyze the ghost noise and the train-test discrepancy in Ghost Batch Normalization. The strength of both effects increases with smaller batch sizes, creating a trade-off between the regularization from the ghost noise and the detrimental effect of train-test discrepancy. We propose a new method, Ghost Noise Injection (GNI), which decouples the ghost noise from the normalization process, allowing us to increase the regularization strength without the need to perform normalization using small batch sizes thereby avoiding their negative effects. It also enables us to apply it in new settings, such as with Layer Normalization, that do not induce noise on its own.
Experimentally, we find that Ghost Nose Injection can provide a greater regularization effect than Ghost Batch Normalization. We ablate the method and find that both the scaling and shifting noise of GNI contributes to its effectiveness. In convolutional neural networks, the noise magnitude can vary between channels and layers depending on their distribution, in particular how much of their variance comes from the spatial extent of a single sample compared to the variance between samples. This “adaptivity” may be important, as we find that simpler IID noise methods are unable to match the regularization effect of GNI. This includes several dropout variants as well as IID noise based on our analysis of the ghost noise distribution.
To summarize, our contributions are as follows:
  • We study the batch size dependency of the regularization effect of batch normalization. We show that both the shift and scale components of the induced noise can positively contribute to the generalization performance, and that the distribution of noise is channel and layer dependent in convolutional networks.
  • We propose a novel regularization method, Ghost Noise Injection (GNI) that decouples the normalization and noise effects of Batch Normalization.
  • We perform extensive experiments to validate the effectiveness of our proposed regularizer, showing that it can outperform Ghost Batch Normalization as well as simpler IID noise methods, such as dropout variants.
@@ 2 Methods and Analysis
2.1 Ghost batch normalization
We can divide a batch ${\bm{X}}\in\mathbb{R}^{B\times C}$ it into $g=\frac{B}{N}$ ghost batches (i.e. non-overlapping subgroups): $\{{\bm{X}}_{1},{\bm{X}}_{2},...{\bm{X}}_{g}\}$, of size $N$ each. Ghost Batch Normalization is equivalent to performing standard Batch Normalization on each ghost batch independently. For simplicity, we will drop the full convolutional tensor notation from this point onward focusing on the simpler fully connected case. We emphasize that the normalization statistics are still computed on a per-feature basis. Writing a ghost batch as ${\bm{X}}_{i}=[{\bm{x}}_{1},{\bm{x}}_{2},..,{\bm{x}}_{N}]\in\mathbb{R}^{B\times C}$, and ${\bm{x}}_{i}\in\mathbb{R}^{1\times C}$ the normalization becomes:
$\widetilde{{\bm{X}}}_{i}=\frac{{\bm{X}}_{i}-{\bm{\mu}}_{i}}{{\bm{\sigma}}_{i}},\qquad\quad{\bm{\mu}}_{i}=\frac{1}{N}\sum_{j}{\bm{x}}_{j}\in\mathbb{R}^{1\times C},\qquad{\bm{\sigma}}^{2}_{i}=\frac{1}{N}\sum_{j}({\bm{x}}_{j}-{\bm{\mu}}_{i})^{2}\in\mathbb{R}^{1\times C}$ (2)
The computation over smaller batch sizes leads to higher variance in ${\bm{\mu}},{\bm{\sigma}}$. Here we note that the term “batch” is highly overloaded and there are at least three different batch sizes of interest:
  • Accelerator batch size is the number of samples each worker (e.g. GPU) uses during a single forward / backward pass through the network.
  • Ghost batch size, or normalization batch size, is the number of samples over which normalization statistics are calculated.
  • Optimization batch size is the number of samples contributing to each optimizer update.
With local accumulation or distributed training, we can have optimization batch size bigger than the accelerator batch size. Likewise, we can also have normalization batch size bigger than accelerator batch size when using e.g. synchronized batch normalization in distributed training.
2.2 Reducing train-test discrepancy
Smaller normalization batch sizes result in increased stochasticity with a potential regularization impact. At the same time, these smaller batches increase the contribution of each sample to the ${\bm{\mu}},{\bm{\sigma}}$ used in the normalization. This does not match the test time behavior, where pre-calculated running statistics are used to normalize new coming samples. To reduce this over-dependency, we explore an alternative where we prevent each sample from contributing to the normalization statistics used for itself. Instead we compute a different set of statistics for each sample, where we consider all other elements in the ghost batch. We refer to this technique as Exclusive Batch Normalization (XBN). This way, the normalization statistics during training no longer depend on the sample itself, better matching the situation at test-time. We can write the XBN mean and variance computation as:
${\bm{\mu}}_{i}=\frac{1}{N-1}\sum_{j\neq i}{\bm{x}}_{j},\qquad{\bm{\sigma}}^{2}_{i}=\frac{1}{N-1}\sum_{j\neq i}({\bm{x}}_{j}-{\bm{\mu}}_{i})^{2}$ (3)

### v2

@@ 1 Introduction
The adoption of normalization techniques has been a prevailing trend in the field of deep learning (Ioffe and Szegedy 2015; Ba, Kiros, and Hinton 2016; Wu and He 2018), notably following the seminal introduction of Batch Normalization (BN)  (Ioffe and Szegedy 2015) in 2015. Batch Normalization operates on each feature independently, normalizing it by subtracting its mean and dividing by its standard deviation. The mean $\bm{\mu}$} and standard deviation (std) $\bm{\sigma}$} with elements corresponding to the different features in the input tensor. During training, these statistics are computed over the batch dimension, causing the network output for a given sample to depend on the others in the batch. This cross-sample dependency is undesirable at inference time, as data can arrive in small correlated batches or even one sample at a time. Consequently, we utilize population-level statistics during inference, instead of computing the mean and variance for individual batches. The population statistics are typically approximated by an exponential moving average of the mean and variance values observed during the training phase.
Figure 1: Impact of ghost batch size on ResNet-18 accuracy on CIFAR-100. Mean (solid) and std (shaded) for 3 runs.
The use of different forms of statistics in training and inference can lead to a detrimental train-test discrepancy resulting in lower test performance. One major difference is that during training each sample contributes to its normalization statistics, biasing the result. The smaller the batch size, the greater the impact of each sample on the statistics, thus a larger bias. We refer to this effect as the self-dependency of batch normalization. The population statistics used during inference are identical for all samples and therefore not biased in the same way. The distribution of normalized activations during testing can therefore be significantly different than during training, leading to worse generalization performance. This effect is studied by Singh and Shrivastava (2019) and Summers and Dinneen (2020) which propose modifying inference to reduce the train-test discrepancy. The use of correlated samples in a batch can similarly lead to a bias causing a different form of train-test discrepancy (Wu and Johnson 2021).
@@
Several alternative methods such as Layer Normalization (Ba, Kiros, and Hinton 2016), Instance Normalization (Huang and Belongie 2017), Group Normalization (Wu and He 2018), Online Normalization (Chiley et al. 2019), Batch Renormalization (Ioffe 2017) and Weight Normalization (Salimans and Kingma 2016) have been proposed to overcome some of the issues with Batch Normalization. These methods either normalize across other dimensions of the tensor $\bm{X}$, apply normalization to the weights instead of activations, or make use of the history over prior batches to augment the effective size of the batch.
Aside from the train-test discrepancy, the dependency of the batch normalization statistics on the elements in the batch can also have a beneficial regularization effect, which is frequently observed but somewhat poorly understood. From a geometrical point of view, Keskar et al. (2017) empirically found that training with small batches arrives at flatter minima, which they argue contributes to better generalization. Hoffer, Hubara, and Soudry (2017) proposed Ghost Batch Normalization (GBN), followed by observations where performing batch normalization over smaller subgroups of the batch could improve generalization. These smaller batches were termed ghost batches by them. Summers and Dinneen (2020) also observed this effect and recommended the use of GBN. The regularization effect arises from the noise in ${\bm{\mu}},{\bm{\sigma}}$ computed over a batch (typically randomly sampled) compared to the corresponding statistics for the full dataset. This noise, which we will refer to as Ghost Noise, arises from the cross-sample dependency and increases for smaller batches, potentially explaining the improved generalization with GBN. Figure 1 shows an example of how the ghost batch size can affect performance.
In this work, we analyze the effects of ghost noise and train-test discrepancy in BN. The strength of both effects increases with smaller batch sizes, creating a trade-off between the regularization from the ghost noise and the detrimental effect of train-test discrepancy. Section 3 constitutes the main part of our analysis and proposed methods. We first propose Exclusive Batch Normalization (XBN). This method mitigates the train-test discrepancy in GBN by simply excluding a sample from the computation of its normalization statistics. This way, self-dependency is eliminated during training, as in inference time. However, it requires moderately sized batches for stable training. Following that, we introduce a novel technique named Ghost Noise Injection (GNI). GNI decouples ghost noise from the normalization process, allowing increased regularization from ghost noise without actually performing normalization using small batches. This method avoids the detrimental effects associated with small batches and can be extended to BN-free settings, such as Layer Normalization, where inherent noise is absent.
Experimentally, we find that GNI can deliver a stronger regularization effect than GBN, resulting in improved test performance across a wide range of training settings. Our ablation study suggests that both the noise in $\bm{\mu}$} (shifting noise) and in $\bm{\sigma}$} (scaling noise) contribute to the regularization effect of GNI. In Convolutional Neural Networks (CNNs), the noise magnitude can vary between channels and layers depending on their distribution, in particular, how much of their variance comes from the spatial extent of a single sample compared to the variance between samples. This “adaptivity” may be important, as we find that simpler i.i.d. (independent and identically distributed) noise methods are unable to match the regularization effect of GNI. This includes several dropout variants as well as i.i.d. noise based on our analysis of the ghost noise distribution.
Summary of Contributions:
  • We examine the impact of smaller batch sizes in GBN, focusing on the potentially useful regularization effect from ghost noise and the harmful train-test discrepancy, offering insights into the trade-off between the two.
  • We introduce Exclusive Batch Normalization (XBN) to mitigate the train-test discrepancy. Our findings also reveal that inherent self-dependency contributes to stability during small-batch training.
  • By deconstructing GBN into two constituent components, we devise a novel method termed Ghost Noise Injection (GNI). This technique allows us to specifically investigate the noise effects associated with small batch sizes.
  • Extensive experimentation highlights the utility of GNI as an effective regularization method across diverse architectures and training scenarios. This underscores the significance of the intrinsic noise within BN.
2 Ghost Batch Normalization
In this section we provide an overview of Ghost Batch Normalization (GBN) from Hoffer, Hubara, and Soudry (2017), followed by a novel modeling of GBN as a composition of GBN and BN, separating the normalization and noise effects.
2.1 Overview
In short, GBN is performed by applying Batch Normalization on disjoint subsets of a batch, i.e. ghost batches. By intentionally operating on smaller batch sizes, GBN increases the stochasticity in the normalization statistics compared to standard batch normalization. This has been found to give a beneficial regularization effect in certain settings, see e.g. Figure 1. Our primary goal in this work is to explore and isolate the effect of Ghost Batch Normalization.
A batch ${\bm{X}}\in\mathbb{R}^{B\times C\times H\times W}$} can be divided into G=B/N{G=B/N} ghost batches $\{{\bm{X}}_{1},{\bm{X}}_{2},...{\bm{X}}_{G}\}$}, of size $N\times C\times H\times W$ each. The output of GBN is obtained by performing batch normalization on each ghost batch. For ghost batch ${\bm{X}}_{g}$ with elements ${X}_{n,c,h,w}$, this can be expressed as:
$\widetilde{{\bm{X}}}_{g}=\frac{({\bm{X}}_{g}-{\bm{\mu}}_{g})}{\sqrt{{\bm{\sigma}}_{g}^{2}+\varepsilon}}$ (1)
where ${\bm{\mu}}_{g},{\bm{\sigma}}^{2}_{g}\in\mathbb{R}^{1\times C\times 1\times 1}$}, ${\bm{\mu}}_{g}=\frac{1}{NHW}\sum_{n,h,w}{X}_{n,c,h,w}$}, ${\bm{\sigma}}^{2}_{g}=\frac{1}{NHW}\sum_{n,h,w}({X}_{n,c,h,w}-{\bm{\mu}}_{g})^{2}$} and we use broadcasted operations similar to PyTorch (Paszke et al. 2019).
Here we note that the term “batch” is highly overloaded, encompassing different batch sizes of interest. Accelerator batch size (denoted $B$) is the number of samples each worker (e.g. GPU) uses during a single forward / backward pass through the network. Normalization batch size or ghost batch size (denoted $N$) is the number of samples over which normalization statistics are calculated.
2.2 Modeling GBN as Double Normalization
In this section, we present a novel perspective on GBN aimed at separating noise effects from normalization effects. The key insight lies in the fact that applying standard batch normalization before ghost batch normalization preserves the output of GBN. For a batch $\bm{X}$, where batch normalization is denoted as BN and ghost batch normalization as GBN, this relationship can be expressed as:
@@ 2.2 Modeling GBN as Double Normalization
which follows from the fact that normalization is invariant to affine transformations of the inputs.
This decomposition of GBN into two successive normalization operations lets us isolate the differences between standard batch normalization and ghost batch normalization. Ignoring $\varepsilon$, the double normalization can be formulated as:
[…]
where we split the normalized batch into $G$ ghost batches and show the following GBN for the $g$-th subgroup. Following the equivalence of equation 1 and equation 4, we have:
$\frac{{\bm{X}}_{g}-{\bm{\mu}}_{g}}{{\bm{\sigma}}_{g}}=\frac{1}{\hat{{\bm{\sigma}}}_{g}}\left(\frac{\bm{X}_{g}-\bm{\mu}}{{\bm{\sigma}}}-\hat{\bm{\mu}}_{g}\right)$ (5)
We can now write the $\bm{\mu}_{g}$ and ${\bm{\sigma}}_{g}$ of the original $\text{GBN}(\bm{X})$ setup in Equation 1 as:
$\bm{\mu}_{g}=\bm{\mu}+{\bm{\sigma}}\hat{\bm{\mu}}_{g},\qquad{\bm{\sigma}}_{g}={\bm{\sigma}}\hat{{\bm{\sigma}}}_{g}$ (6)
GBN would be equivalent to BN if $\hat{\bm{\mu}}_{g}=0$} and $\hat{{\bm{\sigma}}}_{g}=1$}. This happens when the ghost batches are identical, but generally, they contain different samples resulting in random variations in $\hat{\bm{\mu}}_{g}$ and $\hat{{\bm{\sigma}}}_{g}$. These fluctuations induce noise in two components, additive (or shifting) noise from $\hat{\bm{\mu}}_{g}$ and multiplicative (or scaling) noise from $\hat{{\bm{\sigma}}}_{g}$. The increased train-test discrepancy of GBN arises from the dependency of $\hat{\bm{\mu}}_{g}$ and $\hat{{\bm{\sigma}}}_{g}$ on a specific sample (self-dependency).
@@ 3 Methods and Analysis
Within this section, we introduce two innovative approaches designed to enhance the generalization performance of Deep Neural Networks (DNNs): 1) Exclusive Batch Normalization, which addresses the issue of train-test discrepancy from self-dependency; 2) Ghost Noise Injection, a method that enhances training-time stochasticity by incorporating noise from ghost batches. We provide a detailed and clear analysis of this technique.
3.1 Exclusive Batch Normalization (XBN)
We first consider a straightforward technique in which all elements within a ghost batch are utilized except for the element under consideration, for the calculation of normalization statistics. This aims to mitigate the excessive reliance on a single sample during training. We term this approach Exclusive Batch Normalization (XBN). Note that XBN applies a separate set of normalization statistics for each sample. Consequently, the training-time normalization statistics become less dependent on the specific sample itself, aligning more closely with the conditions at test time. The mean and variance of the $k$-th sample in the $g$-th ghost batch of XBN is computed as:
${\bm{\mu}}_{g,k}$ $=\frac{1}{(N-1)HW}\sum_{n\neq k,h,w}{X}_{n,c,h,w}$ (7)
${\bm{\sigma}}^{2}_{g,k}$ $=\frac{1}{(N-1)HW}\sum_{n\neq k,h,w}({X}_{n,c,h,w}-{\bm{\mu}}_{g,k})^{2}$ (8)